Write the expression as a complex number in standard form:
(-2-4)+(3-6i)

Thank You very much!

Answer by Photonjohn(42)

if you mean (-2-4i)(3-6i),
then
-2(3-6i) -4i(3-6i) =
-6 + 12i -12i +24i^2 =
-6 + 24(-1) =
-6 - 24 =
= -30